if a+b=13,a-b=11 find value of a^2+b^2 i am not being able to find it 

Grade:9

2 Answers

Arun
6 years ago
Dear student
In squaring both the equations
a² + b² + 2ab = 169
&
a² + b² - 2ab = 121
Now add
2(a² + b²) = 290
Hence
a² + b² = 145
Regards
Arun (askIITians forum expert)
Sidhant
6 years ago
In sq both equationa^2+b^2+2ab=169anda^2+b^2-2ab=121Now add2(a^2+b^2) =299ans Thanks Sid margin Student in little angel high sec schoolDelhi India near red fort